OpenStudy (anonymous):

How do I translate .12 into a fraction

OpenStudy (anonymous):

.12 = 1/10 + 2/100=10/100+2/100=12/100=3/25

OpenStudy (anonymous):

1 is the tenths digit, 2 is the hundredths digit